# Derivative-Free Guidance in Continuous and Discrete Diffusion Models with Soft Value-Based Decoding

### v3

@@ 7 Conclusion
We propose a novel inference-time algorithm, SVDD, for optimizing downstream reward functions in pre-trained diffusion models that eliminate the need to construct differentiable proxy models. In future work, we plan to conduct experiments in other domains, such as protein sequence optimization (Gruver et al., 2023; Alamdari et al., 2023; Watson et al., 2023) or controllable 3D molecule generation (Xu et al., 2023).
@@ Acknowledgement
Appendix A Difference Between SVDD and “Standard” SMC-Methods
@@
In this section, we compare our algorithm with the SMC-based methods (Wu et al., 2024; Trippe et al., 2022; Dou and Song, 2024; Phillips et al., 2024; Cardoso et al., 2023) for guidance. While they originally aim to solve a conditioning problem, which is different from reward maximization, they algorithms can be converted to reward maximization. We then elaborate on the differences between SVDD, and them in the context of reward maximization. Notably, our SVDD is an instantiation of nested-IS SMC (Naesseth et al., 2019, Algorithm 5) in the literature on computational statistics, whereas these SMC-based Methods rely on standard sequential Monte Carlo.
A.1 SMC-Based Methods
@@
  • In Wu et al. (2024); Cardoso et al. (2023); Phillips et al. (2024), the proposal distribution is not limited to the pre-trained model. Likewise, in our SVDD, we can select an arbitrary proposal distribution, as discussed in Section C.
@@
A.2 Comparison of SVDD with “Standard” SMC-Based Methods (SSM) for Reward Maximization
@@
Now, consider a scenario where the batch size is small, which often occurs due to memory constraints in large pre-trained diffusion models. In this case, while SSM may exhibit suboptimal performance, SVDD-PM can still achieve high performance by choosing a sufficiently large $M$ (assuming that models of the value functions are not significantly larger than those in pre-trained diffusion models).
@@
Appendix B Comparising against DG (Nisonoff et al., 2024) and DiGress (Vignac et al., 2022) in Discrete Diffusion Models
Our method is closely related to guidance methods used in DG (Nisonoff et al., 2024) and DiGress (Vignac et al., 2022). However, we emphasize that our approach is more general, as it operates on any domain, including continuous spaces including Riemannian spaces and discrete spaces, in a unified manner, making a direct comparison not feasible. With this in mind, we provide a comparison focusing on cases where all domains are discrete.
@@
where $\delta t$ is step size. Asymptotically, this is equivalent to sampling from the optimal policy in Theorem 1, as we will show in Remark 5. However, as (Nisonoff et al., 2024) note, sampling from the optimal policy requires $O(KL)$ computation, where $K$ is the vocabulary size and $L$ is the sequence length, which is computationally expensive for large $K$ and $L$. To address this issue, they propose using a Taylor approximation by computing the gradient once. In contrast, we avoid this computational overhead in a different manner, i.e., through importance sampling and resampling. A detailed empirical comparison between our method and theirs is reserved for future work.
@@
Note that in our molecule generation experiment in Section 6, we use GDSS, which operates in continuous space and differs from DiGress.
Appendix C Extension with Arbitrary Proposal Distribution
@@
Appendix D Soft Q-learning
@@
Appendix E Additional Experimental Details
@@
E.1 Additional Setups for Experiments
@@
  • DPS: We require differentiable models that map images to compressibility. For this task, we have used a standard CNN.
@@
  • DPS: Following the implementation in Lee et al. (2023), we use the same GNN model as the reward model. Note that this model cannot compute derivatives with respect to adjacency matrices.
  • SVDD-MC: We use a Graph Isomorphism Network (GIN) model (Xu et al., 2018) as a value function model.
Enhancers.
  • DPS: Although DPS was originally proposed in the continuous space, we have adapted it for our use by incorporating the gradient of the value function model at each step and representing each sequence as a one-hot encoding vector.
  • SVDD-MC: We have used the Enformer model as the function model.
5’UTRs.
We have used ConvGRU as the reward model (Dey and Salem, 2017).
  • DPS: Although DPS was originally proposed in the continuous space, we have adapted it for our use by incorporating the gradient of the value function model at each step and representing each sequence as a one-hot encoding vector.
  • SVDD-MC: We employed ConvGRU as the value function model.
E.2 Additional Results
We evaluate the performance of the generated samples using histograms in Section 6. Here, we present the top 10 and 20 quantiles of the generated samples in Table 2.
Table 2: Top 10 and 20 quantiles of the generated samples for each algorithm. Higher is better.
Domain Quantile Pre-Train Best-N DPS SVDD-MC SVDD-PM
Images compress 20% -86.2 -63.2 -67.1 - -43.7
10% -78.6 -57.3 -61.2 - -38.8
Images aesthetic 20% 5.875 6.246 5.868 - 6.356
10% 5.984 6.343 5.997 - 6.472
Molecules QED 20% 0.771 0.881 0.802 0.912 0.916
10% 0.812 0.902 0.843 0.925 0.928
Molecules SA 20% 0.750 0.916 0.802 1.0 1.0
10% 0.803 0.941 0.849 1.0 1.0
Enhancers 20% 0.74 3.00 2.68 5.53 6.44
10% 1.41 3.52 3.85 5.75 7.02
5’UTR MRL 20% 0.78 0.97 0.90 1.09 1.33
10% 0.86 1.021 0.93 1.12 1.38
5’UTR Stability 20% -0.63 -0.59 -0.62 -0.52 -0.56
10% -0.61 -0.58 -0.60 -0.51 -0.55
Performance of value function training.
We report the performance of value function learning using Monte Carlo regression as follows. We plot the Pearson correlation on the test dataset.
@@ E.2 Additional Results
Figure 5: Training curve
More generated samples.
We have provided additional generated samples in Figure 6, Figure 7, Figure 8, Figure 9, and Figure 10.

### v4

@@ 8 Conclusion
We propose a novel inference-time algorithm, SVDD, for optimizing downstream reward functions in pre-trained diffusion models that eliminate the need to construct differentiable proxy models. Future works include applications in other domains, such as protein sequence optimization (Gruver et al., 2023; Alamdari et al., 2023; Watson et al., 2023) and 3D molecule generation (Xu et al., 2023).
@@ Acknowledgement
Appendix A Further Related Works
Decoding in autoregressive models with rewards.
Appendix B Difference Between SVDD and “Standard” SMC-Methods
In this section, we compare our algorithm with the SMC-based methods (Wu et al., 2024; Trippe et al., 2022; Dou and Song, 2024; Phillips et al., 2024; Cardoso et al., 2023) for guidance. While they originally aim to solve a conditioning problem, which is different from reward maximization, they algorithms can be converted to reward maximization. First, we will explain this converted algorithm. We then elaborate on the differences between SVDD, and them in the context of reward maximization. Notably, our SVDD is an instantiation of nested-IS SMC (Naesseth et al., 2019, Algorithm 5) in the literature on computational statistics, whereas these SMC-based Methods rely on standard sequential Monte Carlo.
B.1 SMC-Based Methods
@@
  • In SMC, resampling is performed across the entire batch. However, in the algorithm, sampling is done within a single batch. Therefore, the algorithms differ significantly. We will discuss the implications in the next section.
@@
  • In Wu et al. (2024); Cardoso et al. (2023); Phillips et al. (2024), the proposal distribution is not limited to the pre-trained model. Likewise, in our SVDD, we can select an arbitrary proposal distribution, as discussed in Section D.
@@
B.2 Comparison of SVDD with “Standard” SMC-Based Methods (SSM) for Reward Maximization
@@
Tailored to optimization in SVDD.
SVDDis considered more suitable for optimization than SMC. This is because, when using SMC for reward maximization, we must set $\alpha$ very low, leading to a lack of diversity. This is expected, as when $\alpha$ approaches 0, the effective sample size reduces to 1. This effect is also evident in our image experiments, as shown in Figure 4. Although SMC performs well in terms of reward functions, there is a significant loss of diversity. Some readers might think this could be mitigated by calculating the effective sample size based on weights (i.e., value functions) and resampling when the effective size decreases; however, this is not the case, as the effective sample size does not directly translate into greater diversity in the generated samples. In contrast, SVDD, maintains much higher diversity.
(a) Samples from SMC
(b) Samples from SVDD-PM
(c) Aesthetic scores from SMC
(d) Aesthetic scores from SVDD-PM
Figure 4: Examples of generated samples from SMC (left) and SVDD (right), when the prompt is “cat.” The histogram of generated samples in terms of the reward function is shown below. Here, the pre-trained models are based on stable diffusion, and we optimize for aesthetic scores. In SMC, we set the batch size $N=60$, and in SVDD, we set the duplication size $M=20$. It is observed that most samples generated by SMC are similar, although diversity in terms of reward functions is roughly maintained. This suggests that the effective sample size in terms of value functions (i.e., weights) does not directly translate to real diversity in the generated samples. On the other hand, in SVDD, the generated samples are much more diverse while still achieving high reward functions.
@@
Now, consider a scenario where the batch size is small, which often occurs due to memory constraints in large pre-trained diffusion models. In this case, while SSM may exhibit suboptimal performance, SVDD-PM can still achieve high performance by choosing a sufficiently large $M$.
@@
Appendix C Comparison against DG (Nisonoff et al., 2024) and DiGress (Vignac et al., 2022) in Discrete Diffusion Models
Our method is closely related to guidance methods used in DG (Nisonoff et al., 2024) and DiGress (Vignac et al., 2022). However, we emphasize that our approach is more general, as it operates on any domain, including continuous spaces including Riemannian spaces and discrete spaces, in a unified manner. In this sense, a strict comparison is not feasible. With this in mind, we provide a comparison focusing on cases where all domains are discrete.
@@
where $\delta t$ is step size. Asymptotically, this is equivalent to sampling from the optimal policy in Theorem 1, as we will show in Remark 3. However, as Nisonoff et al. (2024) note, sampling from the optimal policy requires $O(KL)$ computation, where $K$ is the vocabulary size and $L$ is the sequence length, which is computationally expensive for large $K$ and $L$. To address this issue, they propose using a Taylor approximation by computing the gradient once. However, this is a heuristic in the sense that there is no theoretical guarantee for this approximation. In contrast, we avoid this computational overhead in a different manner, i.e., through importance sampling and resampling. Our algorithm has an asymptotic guarantee as $M$ goes to infinity. Empirically, we have compared two methods in Section 7.
@@
Note that in our molecule generation experiment in Section 7, we use GDSS (Jo et al., 2022), which operates in continuous space and differs from DiGress.
Appendix D Extension with Arbitrary Proposal Distribution
@@
Appendix E Soft Q-learning
@@
Appendix F Additional Experimental Details
@@
F.1 Additional Setups for Experiments
F.1.1 Settings
@@
We define compressibility score as the negative file size in kilobytes (kb) of the image after JPEG compression following (Black et al., 2023). We define aesthetic scorer implemented as a linear MLP on top of the CLIP embeddings, which is trained on more than 400k human evaluations. As pre-trained models, we use Stable Diffusion, which is a common text-to-image diffusion model. As prompts to condition, we use animal prompts following (Black et al., 2023) such as [Dog, Cat, Panda, Rabbit, Horse,…].
@@
We calculate QED and SA scores using the RDKit (Landrum et al., 2016) library. We use the docking program QuickVina 2 (Alhossary et al., 2015) to compute the docking scores following Yang et al. (2021), with exhaustiveness as 1. Note that the docking scores are initially negative values, while we reverse it to be positive and then clip the values to be above 0, i.e.. We compute DS regarding four proteins, parp1 (Poly [ADP-ribose] polymerase-1), 5ht1b (5-hydroxytryptamine receptor 1B), braf (Serine/threonine-protein kinase B-raf), and jak2 (Tyrosine-protein kinase JAK2), which are target proteins that have the highest AUROC scores of protein-ligand binding affinities for DUD-E ligands approximated with AutoDock Vina.
DNA, RNA sequences.
We examine two publicly available large datasets: enhancers ($n\approx 700k$) (Gosai et al., 2023) and UTRs ($n\approx 300k$) (Sample et al., 2019), with activity levels measured by massively parallel reporter assays (MPRA) (Inoue et al., 2019). These datasets have been widely used for sequence optimization in DNA and RNA engineering, particularly in advancing cell and RNA therapies (Castillo-Hair and Seelig, 2021; Lal et al., 2024; Ferreira DaSilva et al., 2024; Uehara et al., 2024). In the Enhancers dataset, each $x$ is a DNA sequence of length 200, while $y\in\mathbb{R}$ is the measured activity in the Hep cell line. In the 5’UTRs dataset, $x$ is a 5’UTR RNA sequence of length 50, and $y\in\mathbb{R}$ is the mean ribosomal load (MRL) measured by polysome profiling.
F.1.2 Baselines and Proposals
We will explain in more detail how to implement baselines and our proposal. We use A100 GPUs for all the tasks.
SVDD-MC.
In SVDD-MC, we require value function models. For images, we use standard CNNs for this purpose, with the same architecture as the reward model. For molecular tasks, we use a Graph Isomorphism Network (GIN) model (Xu et al., 2018) as the value function model. Notably, this model is not differentiable w.r.t. inputs. For GIN, we use mean global pooling and the RELU activation function, and the dimension of the hidden layer is 300. The number of convolutional layers in the GIN model is selected from the set {3, 5}; and we select the maximum number of iterations from {300, 500, 1000}, the initial learning rate from {1e-3, 3e-3, 5e-3, 1e-4}, and the batch size from {32, 64, 128}. For the Enhancer task, we use the Enformer model (Avsec et al., 2021) as the value function model. The Enformer trunk has 7 convolutional layers, each having 1536 channels. as well as 11 transformer layers, with 8 attention heads and a key length of 64. Dropout regularization is applied across the attention mechanism, with an attention dropout rate of 0.05, positional dropout of 0.01, and feedforward dropout of 0.4. The convolutional head for final prediction has 2*1536 input channels and uses average pooling, without an activation function. The model is trained using a batch size selected from {32, 64, 128, 256}, the learning rate from {1e-4, 5e-4, 1e-3}, and the maximum number of iterations from {5k, 10k, 20k}. For the 5’UTR task, we adopt the ConvGRU model (Dey and Salem, 2017). The ConvGRU trunk has a stem input with 4 channels and a convolutional stem that outputs 64 channels using a kernel size of 15. The model contains 6 convolutional layers, each initialized with 64 channels and a kernel size of 5. The convolutional layers use ReLU as the activation function, and a residual connection is applied across layers. Batch normalization is applied to both the convolutional and GRU layers. A single GRU layer with dropout of 0.1 is added after the convolutional layers. The convolutional head for final prediction uses 64 input channels and average pooling, without batch normalization. For training, the batch size is selected from {16, 32, 64, 128}, the learning rate from {1e-4, 2e-4, 5e-4}, and the maximum number of iterations from {2k, 5k, 10k}. All function models are trained to converge in the learning process using MSE loss.
SVDD-PM.
For this proposal, we directly use the reward feedback to evaluate. We remark when the reward feedback is also learned from offline data, technically, it would be better to use techniques mitigating over-optimization as discussed in Uehara et al. (2024). However, since this point is tangential in our work, we don’t do it.
DPS.
We require differentiable models. For this task, for images, enhancers, and 5’UTRs, we use the same method as SVDD-MC. For molecules, we follow the implementation in Lee et al. (2023), and we use the same GNN model as the reward model. Note that we cannot compute derivatives with respect to adjacency matrices when using the GNN model. Regarding $\alpha$, we choose several candidates and report the best one. For image tasks we select from [5.0, 10.0] and for bio-sequence tasks we select from [1.0, 2.0]. For molecule QED task we select from {0.2, 0.3, 0.4, 0.5}, for molecule SA task {0.1, 0.2, 0.3}, and for molecule docking tasks we select from {0.4, 0.5, 0.6}.
SMC.
For value function models, we use the same method as SVDD-PM. Regarding $\alpha$, we choose several candidates and report the best one. For image tasks we select from [10.0, 40.0]. For Enhancer and 5’UTR tasks as well as molecule QED and SA tasks we select from {0.1, 0.2, 0.3, 0.4}, while for molecule docking tasks we select from {1.5, 2.0, 2.5}.
F.2 Software and Hardware
Our implementation is under the architecture of PyTorch (Paszke et al., 2019). The deployment environments are Ubuntu 20.04 with 48 Intel(R) Xeon(R) Silver, 4214R CPU @ 2.40GHz, 755GB RAM, and graphics cards NVIDIA RTX 2080Ti. Each of our experiments is conducted on a single NVIDIA RTX 2080Ti or RTX A6000 GPU.
F.3 Additional Results
Histograms.
In the main text, we present several quantiles. Here, we plot the reward score distributions of generated samples as histograms in Figure 5.
@@ F.3 Additional Results
(d) Molecules: SA
(e) Molecules: DS - parp1
(f) Molecules: DS - 5ht1b
(g) Molecules: DS - jak2
(h) Molecules: DS - braf
(i) Enhancers
(j) 5’UTRs: MRL
Figure 5: We show the histogram of generated samples in terms of reward functions. We consistently observe that SVDD demonstrates strong performances.
Performance of value function training.
We report the performance of value function learning using Monte Carlo regression as follows in SVDD-MC. In Figure 6, we plot the Pearson correlation on the test dataset for the Enhancer and 5’UTR tasks, as well as the test MSE for the molecular task of parp1 docking score.
(a) 5’UTR MRL
(b) Enhancers
(c) Molecules
Figure 6: Training curve of value functions
Validity metrics for molecule generation.
To evaluate the validity of our method in molecule generation, we report several key metrics that capture different aspects of molecule quality and diversity in Table 3 on page 3.
Table 3: Comparison of the generated molecules of pre-trained GDSS model and SVDD applied on various metrics.
Method Valid Unique Novelty FCD SNN Frag/Test Scaf NSPDK MMD Mol Stable Atm Stable
Pre-trained 1.0 1.0 1.0 22.2799 0.2992 0.8274 0.0033 0.0260 0.2903 0.9256
SVDD 1.0 0.9375 1.0 21.5671 0.3441 0.7803 0.0838 0.0772 0.4783 0.9095
The validity of a molecule indicates its adherence to chemical rules, defined by whether it can be successfully converted to SMILES strings by RDKit. Uniqueness refers to the proportion of generated molecules that are distinct by SMILES string. Novelty measures the percentage of the generated molecules that are not present in the training set. Fréchet ChemNet Distance (FCD) measures the similarity between the generated molecules and the test set. The Similarity to Nearest Neighbors (SNN) metric evaluates how similar the generated molecules are to their nearest neighbors in the test set. Fragment similarity measures the similarity of molecular fragments between generated molecules and the test set. Scaffold similarity assesses the resemblance of the molecular scaffolds in the generated set to those in the test set. The neighborhood subgraph pairwise distance kernel Maximum Mean Discrepancy (NSPDK MMD) quantifies the difference in the distribution of graph substructures between generated molecules and the test set considering node and edge features. Atom stability measures the percentage of atoms with correct bond valencies. Molecule stability measures the fraction of generated molecules that are chemically stable, i.e., whose all atoms have correct bond valencies. Specifically, atom and molecule stability are calculated using conformers generated by RDKit and optimized with UFF (Universal Force Field) and MMFF (Merck Molecular Force Field).
We compare the metrics using 512 molecules generated from the pre-trained GDSS model and from SVDD optimizing SA, as shown in Table 3. Overall, our method achieves comparable performances with the pre-trained model on all metrics, maintaining high validity, novelty, and uniqueness while outperforming on several metrics such as molecule stability, FCD, SNN, and scaffold similarity. These results indicate that our approach can generate a diverse set of novel molecules that are chemically plausible and relevant.
More Ablation Studies.
We provide several more ablation studies regarding $M$ on top of the results in the main text, as plotted in Figure 7. The results are consistent with what we have observed in Figure 3.
(a) Performance of SVDD as $M$ varies.
(b) GPU time and max memory of SVDD as $M$ varies.
Figure 7: Abltation Studies (for image generation while optimizing the compressibility)
Visualization of more generated samples.
We provide additional generated samples in this section. Figure 8 and Figure 9 show comparisons of generated images from baseline methods and SVDD with different M values regarding compressibility and aesthetic score, respectively. Figure 10 and Figure 11 presents the comparisons of visualized molecules generated from the baseline model and SVDD regarding QED and SA, respectively. The visualizations validate the strong performances of SVDD. Given that SVDD can achieve optimal SA for many molecules, we also visualize some molecules with optimal SA generated by SVDD, as shown in Figure 12. In Figure 13, Figure 14, Figure 15, and Figure 16 we visualizes the docking of SVDD-generated molecular ligands to proteins parp1, 5ht1b, jak2, and braf, respectively. Docking scores presented above each column quantify the binding affinity of the ligand-protein interaction, while the figures include various representations and perspectives of the ligand-protein complexes. We aim to provide a complete picture of how each ligand is situated within both the local binding environment and the larger structural framework of the protein. First rows show close-up views of the ligand bound to the protein surface, displaying the topography and electrostatic properties of the protein’s binding pocket and providing insight into the complementarity between the ligand and the pocket’s surface. Second rows display distant views of the protein using the surface representation, offering a broader perspective on the ligand’s spatial orientation within the global protein structure. Third rows provide close-up views of the ligand interaction using a ribbon diagram, which represents the protein’s secondary structure, such as alpha-helices and beta-sheets, to highlight the specific regions of the protein involved in binding. Fourth rows show distant views of the entire protein structure in ribbon diagram, with ligands displayed within the context of the protein’s full tertiary structure. Ligands generally fit snugly within the protein pocket, as evidenced by the close-up views in both the surface and ribbon diagrams, which show minimal steric clashes and strong surface complementarity.
@@
Figure 13: Additional generated samples from SVDD (Domain: Molecules, Reward: Docking score - parp1 (normalized as $max(-DS,0)$))
Figure 14: Additional generated samples from SVDD (Domain: Molecules, Reward: Docking score - 5ht1b (normalized as $max(-DS,0)$))
Figure 15: Additional generated samples from SVDD (Domain: Molecules, Reward: Docking score - jak2 (normalized as $max(-DS,0)$))
Figure 16: Additional generated samples from SVDD (Domain: Molecules, Reward: Docking score - braf (normalized as $max(-DS,0)$))